# Infinite-Dimensional Stochastic Differential Equations and Diffusion Dynamics of Coulomb Random Point Fields

### v3

@@ 5.1. Construction of the logarithmic derivative $\mathsf{d}_{R,\mathsf{y}}$ of $\mu_{R,\mathsf{y}}$
The convergence in (5.38) holds in $C^{1}(\overline{S}_{R,\epsilon}(\mathsf{s}))$ for each $\epsilon>0$ and all $\mathsf{s}$ such that $\mathsf{s}=\sum_{s^{i}\in\overline{S}_{R}}\delta_{s^{i}}$, where $\overline{S}_{R,\epsilon}(\mathsf{s})=\{{x}\in\overline{S}_{R};\lvert{x}-s^{i}\rvert\geq\epsilon\text{ for all }i\}$.
From Lemma 5.5 and (5.38), we obtain weakly in $L^{1}(\Lambda_{R}^{[1]})$
@@
$=\lim_{\mathfrak{n}\to\infty}\Big{(}\beta\Big{(}\sum_{\mathbf{i}\in\mathbb{I}(\ell)}\eta_{R,\mathsf{y}}^{\mathbf{i},\mathfrak{n},\mathfrak{n}}{x}^{\mathbf{i}}\Big{)}-\mathsf{d}_{R,\mathsf{y}^{N_{\mathfrak{n}}}}^{N_{\mathfrak{n}}}({x},\mathsf{s})+\mathsf{d}_{R,\mathsf{y}^{N_{\mathfrak{n}}}}^{N_{\mathfrak{n}}}({x},\mathsf{s})\Big{)}\mathsf{f}_{Q}\mathfrak{m}_{R,\mathsf{y}^{N_{\mathfrak{n}}}}^{N_{\mathfrak{n}},[1]}$
$=\Big{(}\beta\Big{(}-\mathscr{R}_{R,\mathsf{y}}^{\ell}({x})+\nabla\Phi({x})+\sum_{s^{i}\in\overline{S}_{R}}\nabla_{{x}}\Psi({x},s^{i})\Big{)}+\mathsf{d}_{R,\mathsf{y}}({x},\mathsf{s})\Big{)}\mathsf{f}_{Q}\mathfrak{m}_{R,\mathsf{y}}^{[1]}.$ (5.39)
Here, we used (5.14) and (5.38) for the last line.
From $\mu_{R,\mathsf{y}}^{[1]}(\overline{S}_{R}\times\overline{\mathsf{S}}_{R})>0$, we have $\int_{\overline{S}_{R}\times\overline{\mathsf{S}}_{R}}\mathfrak{m}_{R,\mathsf{y}}^{[1]}d\Lambda_{R}^{[1]}>0$. Hence, the elements $\{\beta{x}^{\mathbf{i}}\mathsf{f}_{Q}\mathfrak{m}_{R,\mathsf{y}}^{[1]};\mathbf{i}\in\mathbb{I}(\ell)\}$ are linearly independent in $L^{1}(\Lambda_{R}^{[1]})$. From this and (5.39), each numerical sequence $\{\eta_{R,\mathsf{y}}^{\mathbf{i},\mathfrak{n},\mathfrak{n}}\}_{\mathfrak{n}\in\mathbb{N}}$ has a finite limit $\eta_{R,\mathsf{y}}^{\mathbf{i}}$ such that
$\lim_{\mathfrak{n}\to\infty}\sum_{\mathbf{i}\in\mathbb{I}(\ell)}\eta_{R,\mathsf{y}}^{\mathbf{i},\mathfrak{n},\mathfrak{n}}\big{(}\beta{x}^{\mathbf{i}}\mathsf{f}_{Q}\mathfrak{m}_{R,\mathsf{y}}^{[1]}\big{)}=\sum_{\mathbf{i}\in\mathbb{I}(\ell)}\eta_{R,\mathsf{y}}^{\mathbf{i}}\big{(}\beta{x}^{\mathbf{i}}\mathsf{f}_{Q}\mathfrak{m}_{R,\mathsf{y}}^{[1]}\big{)},$
which implies the first equality in (5.35), that is, $\lim_{\mathfrak{n}\to\infty}\eta_{R,\mathsf{y}}^{\mathbf{i},\mathfrak{n},\mathfrak{n}}=\eta_{R,\mathsf{y}}^{\mathbf{i}}$ for each $\mathbf{i}\in\mathbb{I}(\ell)$.
Let $\mathscr{X}_{R}(\ell)$ and ${x}_{\mathbf{j}}^{\mathbf{i}}$ for ${x}\in\mathscr{X}_{R}(\ell)$ be as in (2.25). Then, for each $\mathbf{j}\in\mathbb{I}(\ell)$,

### v4

@@ 10.1. Tail decomposition of Dirichlet forms and weak solution
$\{\mathbf{S}_{\mu_{\mathsf{a}}}\}$ is also clear because the $m$-labeled process $\mathbf{X}^{[m]}$ given by $\mathbf{X}$ is associated with a symmetric Dirichlet form on $L^{2}(\mu_{\mathsf{a}}^{[m]})$.
@@
Let $\mathbf{M}=(M^{i})_{i=1}^{\infty}$ such that $M_{t}^{i}=\int_{0}^{t}\sigma(X_{u}^{i})dB_{u}^{i}$. Because $\mathfrak{a}=\sigma^{t}\sigma$, $\sigma\in C_{b}^{1}(\mathbb{R}^{d})$, and $\mathfrak{a}$ is uniformly elliptic, we can represent $\mathbf{B}=(B^{i})_{i=1}^{\infty}$ as a function of $\mathbf{M}$. Thus, $\mathbf{B}$ is a $\sigma[\mathbf{M}_{s};s\leq t]$-adapted process. Because $(\mathbf{X},\mathbf{B})$ is a solution of ISDE (2.33),